# Searching for light long-lived particles at SHiNESS

### v2

@@ Abstract
Recently Soleti et al. [JHEP03(2024)148] proposed a new experiment called SHiNESS at the upcoming European Spallation Source (ESS) facility, making use of the 2-GeV proton beam there impinging on a fixed target, in order to search for hidden sterile neutrinos that could lie in different mass ranges and arise with distinct signatures. Such signatures include excesses in electron-positron pairs that may originate from displaced decays of long-lived particles (LLPs). At the ESS, the dominant sources of such LLPs are decays at rest of $\pi^{+}$ mesons and $\mu^{+}$ leptons. We choose to investigate theoretical scenarios of long-lived light binos in the R-parity-violating supersymmetry and long-lived weak-violating electrophilic axionlike particles, as these LLPs can be produced from decays of $\pi^{+}$ and $\mu^{+}$. Since the $\pi^{+}$’s and $\mu^{+}$’s decay at rest at the ESS, we compute the spectra of the therefrom produced LLPs, and thus estimate the expected sensitivity reach of SHiNESS to the LLPs in these two scenarios. Our calculation shows that in most of the relevant benchmark scenarios, SHiNESS can probe large parameter regions of these models beyond the existing bounds, in just a couple of years of data-collection time.
@@ 1 Introduction
Concretely, we consider light binos in the R-parity-violating supersymmetry (RPV-SUSY) Barbier:2004ez ; Dreiner:1997uz ; Mohapatra:2015fua , and weak-violating electrophilic axionlike particles (ALPs) Altmannshofer:2022ckw , as benchmark models. In the RPV-SUSY, the lightest neutralino with a GeV-scale mass Domingo:2022emr , which is necessarily bino-like Gogoladze:2002xp ; Dreiner:2009ic , is still allowed by all astrophysical and cosmological bounds Grifols:1988fw ; Ellis:1988aa ; Lau:1993vf ; Dreiner:2003wh ; Dreiner:2013tja ; Profumo:2008yg ; Dreiner:2011fp if we lift the GUT relation between the gauginos, $M_{1}\approx 0.5\,M_{2}$ Choudhury:1995pj ; Choudhury:1999tn , the light bino is unstable to avoid overclosing the Universe Bechtle:2015nua , and the dark matter does not consist of the lightest neutralino Belanger:2002nr ; Hooper:2002nq ; Bottino:2002ry ; Belanger:2003wb ; AlbornozVasquez:2010nkq ; Calibbi:2013poa . The lightest neutralino should also be the lightest supersymmetric particle. The light bino can decay via RPV couplings, and if its mass is small, it has a long lifetime. Recent phenomenological studies on long-lived light binos in the RPV-SUSY can be found in, e.g. Refs. deVries:2015mfw ; Cottin:2022gmk ; Gehrlein:2021hsk ; Dercks:2018eua ; Dreiner:2020qbi ; Dreiner:2022swd ; Gunther:2023vmz .
@@
ALPs could solve the strong CP problem ParticleDataGroup:2022pth ; Peccei:1977ur ; Peccei:2006as and in addition, serve as a dark-matter candidate Dine:1982ah ; Abbott:1982af ; Preskill:1982cy ; Marsh:2015xka ; Lambiase:2018lhs ; Auriol:2018ovo ; Houston:2018vrf , among other motivations. ALPs are often conceived as arising from the breaking of a global U(1) Peccei-Quinn (PQ) Peccei:1977hh symmetry as pseudo-Nambu-Goldstone bosons. They are predicted to be long-lived for their feeble couplings with SM particles, even with masses in the MeV-GeV range. In this work, we focus on electrophilic ALPs with weak-violating couplings Altmannshofer:2022ckw . In this case, a four-point vertex arises between an ALP, a $W$-boson, an electron, and an electron neutrino, with no helicity suppression present. Some recent phenomenological works covering long-lived ALPs include Refs. Bertholet:2021hjl ; Co:2022bqq ; Carmona:2022jid ; Cheung:2022umw ; Cheung:2021mol ; Cheung:2024qve ; Lu:2022zbe .
@@
We emphasize that for the installation of the SHiNESS experiment, no upgrade to the beamline is required, and the experiment can start taking data as soon as the ESS launches its operation. Moreover, as we will show, strong limits beyond current bounds on the considered models can be obtained in just a few years, strongly motivating the implementation of SHiNESS.
@@
This paper is structured as follows. In Sec. 2 we introduce the theoretical models we study, and in Sec. 3 we detail the setup of the SHiNESS experiment. Sec. 4 is devoted to presenting numerical results of the projected sensitivity reach of SHiNESS. At the end, we conclude the work in Sec. 5 with a summary of our findings.
@@ 2.1 Light binos in the RPV-SUSY
where the operators in the first and the second lines do not conserve the lepton number and the baryon number, respectively, and $i,j,k=1,2,3$ are generation indices. $\kappa_{i}$ are couplings of mass dimension 1, and $\lambda$, $\lambda^{\prime}$, and $\lambda^{\prime\prime}$ are dimensionless couplings. Here, the factor $\frac{1}{2}$ is attached to the operators $LL\bar{E}$ and $\bar{U}\bar{D}\bar{D}$ because these operators possess the anti-symmetry property such that $\lambda_{ijk}=-\lambda_{jik}$ and $\lambda^{\prime\prime}_{ijk}=-\lambda^{\prime\prime}_{ikj}$. In this work, we assume that only $\lambda$ and $\lambda^{\prime}$ couplings with certain flavor combinations are non-zero; this can be theoretically realized, as the bilinear term can be rotated away at a given energy scale Hall:1983id ; Dreiner:2003hw and the BNV term can be vanishing if, for example, the baryon-triality symmetry is imposed Dreiner:2005rd ; Ibanez:1991hv ; Martin:1997ns ; Grossman:1998py ; Dreiner:2006xw ; Dreiner:2007uj ; Bernhardt:2008jz .
@@
As discussed in Sec. 1, a light MeV-scale bino is allowed and can be long-lived for tiny RPV couplings. For the SHiNESS experiment, we focus on LLPs produced in $\pi^{+}$ and $\mu^{+}$ decays and on the final-state signature of an electron-positron pair from the $\tilde{\chi}^{0}_{1}$ decays. The $\pi^{+}$ decays into a light bino can be induced by the $\lambda^{\prime}_{111}$ or $\lambda^{\prime}_{211}$ couplings, while the signal decays of the $\mu^{+}$ lepton and the bino can be mediated by various $\lambda$ couplings. In Table 1 we summarize the possible RPV benchmark scenarios to study. In each benchmark scenario, we assume only two RPV LNV couplings are non-vanishing, mediating production and decay of the lightest neutralino, respectively, except in one benchmark B3.1 where only a single coupling $\lambda_{121}$ is assumed to be non-zero and it leads to both the production and decay of the light bino. Moreover, in certain benchmarks the decay coupling can also induce kinematically allowed production mode of the binos; these contributions are taken into account. Specifically, in Table 1, we list the production and decay couplings for the light bino, as well as the corresponding signal processes. We note that in benchmark scenarios B1.1 and B2.1, $\lambda_{121}$ are taken as the decay coupling. However, in principle, it also induces bino production from $\mu^{+}$-lepton decay, $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$; this production mode is not included in our computation because in these benchmarks we focus on bino production in $\pi^{+}$ decays. On the other hand, in benchmark scenarios B4.1, B5.1, and B6.1, since we are considering binos produced from $\mu^{+}$-lepton decays, we include the contributions of the decay coupling $\lambda_{121}$ to the bino production from $\mu^{+}$ decays. In addition, for kinematical reasons, the production couplings in general induce no decays of the light bino in the interested mass range of the listed benchmark scenarios. One exception is in B1.1 and B1.2, where the decay $\tilde{\chi}^{0}_{1}\to\pi^{0}\overset{(-)}{\nu}_{e}$ can, in principle, take place via $\lambda^{\prime}_{111}$, but the allowed bino mass window, $m_{\pi^{0}}<m_{\tilde{\chi}^{0}_{1}}<m_{\pi^{+}}$, is tiny, and therefore we ignore this case. Also, in B3.1 by definition the production and decay couplings are identical. Finally, in B3.2, the production coupling $\lambda_{121}$ can induce the $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ decays which are taken into account in the numerical computation.
@@
B1.1 $\lambda^{\prime}_{111}$ $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$-
B1.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
B2.1 $\lambda^{\prime}_{211}$ $\pi^{+}\to\mu^{+}\tilde{\chi}^{0}_{1}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$-
B2.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
B3.1 $\lambda_{121}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ Redundant Redundant
B3.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$
B4.1 $\lambda_{122}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\nu_{\mu}e^{+}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$-
B4.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
B5.1 $\lambda_{321}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{\tau}e^{+}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$-
B5.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
B6.1 $\lambda_{312}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\nu_{\tau}e^{+}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$-
B6.2 $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
@@
Table 1: Summary of the benchmark scenarios of the light binos in the RPV-SUSY. For benchmarks B3.1-B6.2, the $\mu^{+}$ leptons are produced in $\pi^{+}\to\mu^{+}\nu_{\mu}$ decays at rest at the ESS. The lightest neutralino is of Majorana nature and can hence decay to both charge-conjugated channels. “--” means either theoretically irrelevant, or kinematically forbidden.
@@
To compute the production rates of the light bino from pion and muon decays via $\lambda^{\prime}$ and $\lambda$ couplings, respectively, we resort to the analytic expressions provided in Refs. deVries:2015mfw ; Dreiner:2023gir , while the decay widths’ computation of the light bino via $\lambda$ couplings follows the computation procedure given in Ref. Dreiner:2023gir .
@@
We now discuss the present bounds on the relevant RPV couplings. These bounds stem from either measurements on low-energy observables Allanach:1999ic or recast of past searches for long-lived HNLs Dreiner:2023gir .
@@
For the coupling $\lambda^{\prime}_{111}$, both measurements of neutrinoless double-beta ($0\nu\beta\beta$) decays and existing bounds on long-lived HNLs provide the leading limits on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the lightest neutralino’s mass $m_{\tilde{\chi}^{0}_{1}}$. The bound from $0\nu\beta\beta$ decays on the coupling is expressed as follows Mohapatra:1986su ; Hirsch:1995zi ; Hirsch:1995ek ; Babu:1995vh ,
$\lambda^{\prime}_{111}<5.2\cdot 10^{-4}\cdot\Big{(}\frac{m_{\tilde{e}}}{100\text{ GeV}}\Big{)}^{2}\cdot\sqrt{\frac{m_{\tilde{\chi}^{0}_{1}}}{100\text{ GeV}}},$ (2)
where $m_{\tilde{e}}$ is the selectron’s mass. The reinterpretation of the HNL searches’ results in terms of the light binos in the RPV-SUSY was performed in Ref. Dreiner:2023gir , and its Fig. 2 shows that these bounds are stronger than those from the $0\nu\beta\beta$ decays for $m_{\tilde{\chi}^{0}_{1}}$ roughly between 30 MeV and 130 MeV.
Similarly, bounds on $\lambda^{\prime}_{211}$ have been obtained from both measurements on $R_{\pi}=\Gamma(\pi\to e\nu)/\Gamma(\pi\to\mu\nu)$ Barger:1989rk and recast of HNL searches Dreiner:2023gir . The former gives a bound of
$\lambda^{\prime}_{211}<0.059\times\frac{m_{\tilde{d}_{R}}}{100\text{ GeV}},$ (3)
where $m_{\tilde{d}_{R}}$ labels the mass of a right-handed sdown quark. However, Fig. 2 of Ref. Dreiner:2023gir shows that reinterpretation of the HNL searches leads to the strongest bounds on $\lambda^{\prime}_{211}/m^{2}_{\tilde{f}}$ for the light bino’s mass below approximately 33 MeV.
The involved $\lambda$ couplings all receive bounds from low-energy processes, as listed in Ref. Allanach:1999ic , and we reproduce them here:
$\lambda_{121}$ $<0.049\cdot\frac{m_{\tilde{e}_{R}}}{100\text{ GeV}},\lambda_{122}<0.049\cdot\frac{m_{\tilde{\mu}_{R}}}{100\text{ GeV}},$ (4)
$\lambda_{131}$ $<0.062\cdot\frac{m_{\tilde{e}_{R}}}{100\text{ GeV}},$ (5)
$\lambda_{321}$ $<0.070\cdot\frac{m_{\tilde{e}_{R}}}{100\text{ GeV}},\lambda_{312}<0.062\cdot\frac{m_{\tilde{\mu}_{R}}}{100\text{ GeV}}.$ (6)
Here, $\tilde{e}_{R}$ and $\tilde{\mu}_{R}$ denote right-handed selectron and smuon, respectively. Among these constraints, the bounds on $\lambda_{121}$ and $\lambda_{122}$ originate from charged-current universality Barger:1989rk . The upper limits on $\lambda_{131}$ were derived from measurements of $R_{\tau}=\Gamma(\tau\to e\nu\bar{\nu})/\Gamma(\tau\to\mu\nu\bar{\nu})$ Barger:1989rk . Moreover, current bounds on $\lambda_{321}$ and $\lambda_{312}$ stem from measurements on $R_{\tau\mu}=\Gamma(\tau\to\mu\nu\bar{\nu})/\Gamma(\mu\to e\nu\bar{\nu})$ Ledroit:1998 .
@@
In addition, the $\lambda_{121}$, $\lambda_{122}$, and $\lambda_{321}$ couplings are also bounded from uncertainty on the muon decay width of $\mu^{\pm}\to e^{\pm}+$invisible ParticleDataGroup:2022pth ($\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$) as functions of the $m_{\tilde{\chi}^{0}_{1}}$, as given in Fig. 6 of Ref. Dreiner:2023gir . The results are comparable with the bounds obtained from low-energy observables.
Besides the single-coupling bounds, there are strong constraints on the product of certain pairs of RPV couplings. Relevant for us are the bounds shown below:
$\sqrt{\lambda_{121}\lambda_{122}}<8.4\times 10^{-4}\frac{m_{\tilde{\nu}_{\mu L}}}{100\text{ GeV}},$ (7)
@@
$\sqrt{\lambda_{121}\lambda^{\prime}_{111}}<2.0\times 10^{-4}\frac{m_{\tilde{\nu}_{eL}}}{100\text{ GeV}},$ (8)
where the two bounds were derived from considerations of measurements on $\mu\to 3e$ Choudhury:1996ia and on $\mu\text{Ti}\to e\text{Ti}$ Kim:1997rr ; Huitu:1997bi , respectively, and $m_{\tilde{\nu}_{\mu L}}$ and $m_{\tilde{\nu}_{eL}}$ are masses of left-handed muon sneutrino and electron sneutrino, respectively.
@@
Finally, in this work, we assume degenerate sfermion masses $m_{\tilde{f}}$ for simplicity.

### v3

# Searching for long-lived light neutralinos and axionlike particles at the SHiNESS experiment

@@ Abstract
Recently Soleti et al. [JHEP03(2024)148] proposed a new experiment called SHiNESS at the upcoming European Spallation Source (ESS) facility, making use of the 2-GeV proton beam there impinging on a fixed target, in order to search for hidden sterile neutrinos that could lie in different mass ranges and arise with distinct signatures. Such signatures include excesses in electron-positron pairs that may originate from displaced decays of long-lived particles (LLPs). At the ESS, the dominant sources of such LLPs are decays at rest of $\pi^{+}$ mesons and $\mu^{+}$ leptons. We choose to investigate theoretical scenarios of long-lived light neutralinos in the R-parity-violating supersymmetry and long-lived weak-violating electrophilic axionlike particles, as these LLPs can be produced from decays of $\pi^{+}$ and $\mu^{+}$. Since the $\pi^{+}$’s and $\mu^{+}$’s decay at rest at the ESS, we compute the spectra of the therefrom produced LLPs, and thus estimate the expected sensitivity reach of SHiNESS to the LLPs in these two scenarios. Our calculation shows that in multiple relevant benchmark scenarios, SHiNESS can probe large parameter regions of these models beyond the existing bounds, in just a couple of years of data-collection time.
@@ 1 Introduction
Concretely, we consider light binos in the R-parity-violating supersymmetry (RPV-SUSY) Barbier:2004ez ; Dreiner:1997uz ; Mohapatra:2015fua , and weak-violating electrophilic axionlike particles (ALPs) Altmannshofer:2022ckw , as benchmark models. In the RPV-SUSY, the lightest neutralino with a GeV-scale or even smaller mass Domingo:2022emr , which is necessarily bino-like Gogoladze:2002xp ; Dreiner:2009ic , is still allowed by all astrophysical and cosmological bounds Grifols:1988fw ; Ellis:1988aa ; Lau:1993vf ; Dreiner:2003wh ; Dreiner:2013tja ; Profumo:2008yg ; Dreiner:2011fp if we lift the GUT relation between the gauginos, $M_{1}\approx 0.5\,M_{2}$ Choudhury:1995pj ; Choudhury:1999tn , the light bino is unstable to avoid overclosing the Universe Bechtle:2015nua , and the dark matter does not consist of the lightest neutralino Belanger:2002nr ; Hooper:2002nq ; Bottino:2002ry ; Belanger:2003wb ; AlbornozVasquez:2010nkq ; Calibbi:2013poa . The lightest neutralino should also be the lightest supersymmetric particle. The light bino can decay via RPV couplings, and if its mass is small, it has a long lifetime. We note that a priori the R-parity can be either conserved or broken, and therefore the RPV-SUSY is equally legitimate as the R-parity-conserving SUSY (RPC-SUSY) Ibanez:1991pr ; Dreiner:2005rd . Moreover, the RPV-SUSY offers a solution to the non-vanishing active-neutrino masses Hall:1983id ; Grossman:1998py ; Hirsch:2000ef ; Dreiner:2006xw ; Dreiner:2011ft and a richer phenomenology at colliders than the RPC-SUSY Dreiner:1991pe ; deCampos:2007bn ; Dercks:2017lfq ; Dreiner:2023bvs . It could also explain several experimental anomalies reported in recent years, including the $B$-anomalies Trifinopoulos:2019lyo ; Hu:2020yvs ; BhupalDev:2021ipu , the ANITA anomaly Collins:2018jpg ; Altmannshofer:2020axr , as well as the muon $g-2$ Hu:2019ahp ; Zheng:2021wnu ; BhupalDev:2021ipu . The light binos in the RPV-SUSY can be produced in rare decays of, e.g. bottom and charm mesons deVries:2015mfw . Particularly, very few studies exist that constrain such long-lived light binos that are produced from decays of pions or muons, cf. for instance, Refs. Choudhury:1999tn ; Dreiner:2023gir . Studying the sensitivity of the SHiNESS experiment to this theoretical scenario thus complements the existing efforts. Recent phenomenological studies on long-lived light binos in the RPV-SUSY can be found in, e.g. Refs. deVries:2015mfw ; Cottin:2022gmk ; Gehrlein:2021hsk ; Dercks:2018eua ; Dreiner:2020qbi ; Dreiner:2022swd ; Gunther:2023vmz .
@@
ALPs could solve the strong CP problem ParticleDataGroup:2022pth ; Peccei:1977ur ; Peccei:2006as and in addition, serve as a dark-matter candidate Dine:1982ah ; Abbott:1982af ; Preskill:1982cy ; Marsh:2015xka ; Lambiase:2018lhs ; Auriol:2018ovo ; Houston:2018vrf , among other motivations. They are often conceived as arising from the breaking of a global U(1) Peccei-Quinn (PQ) Peccei:1977hh symmetry as pseudo-Nambu-Goldstone bosons. They are predicted to be long-lived for their feeble couplings with Standard-Model (SM) particles, even with masses in the MeV-GeV range. The ALPs can couple to SM leptons, quarks, gauge bosons, etc. Among these various possibilities, leptophilic ALPs are motivated for their connections to leptonic electric dipole moment Kirpichnikov:2020lws , charged lepton flavor violation Han:2020dwo ; Cheung:2021mol ; Bertuzzo:2022fcm , explaining the neutrino accesses at MiniBooNE Chang:2021myh , or accommodating the discrepancies observed between the measurements of the leptonic anomalous magnetic moment and the SM predictions Bauer:2019gfk ; Cornella:2019uxs ; Buen-Abad:2021fwq ; Bauer:2021mvw ; Davoudiasl:2024vje . In this work, we focus on electrophilic ALPs with weak-violating couplings Altmannshofer:2022ckw . In this case, a four-point vertex arises between an ALP, a $W$-boson, an electron, and an electron neutrino, with no helicity suppression present. Some recent phenomenological works covering long-lived ALPs include Refs. Bertholet:2021hjl ; Co:2022bqq ; Carmona:2022jid ; Cheung:2022umw ; Cheung:2021mol ; Cheung:2024qve ; Lu:2022zbe ; Buonocore:2023kna .
@@
We emphasize that for the installation of the SHiNESS experiment, no upgrade to the beamline is required, and the experiment can start taking data as soon as the ESS launches its operation. Moreover, as we will show, strong limits beyond current bounds on the considered models can be obtained in just a few years, strongly motivating the implementation of the SHiNESS experiment.
@@
This paper is structured as follows. In Sec. 2 we introduce the theoretical models we study, and in Sec. 3 we detail the setup of the SHiNESS experiment. Sec. 4 is devoted to presenting numerical results of the projected sensitivity reach of SHiNESS. At the end, we summarize our findings in Sec. 5.
@@ 2.1 Light binos in the RPV-SUSY
where the operators in the first and the second lines break the lepton number and the baryon number, respectively, and $i,j,k=1,2,3$ are generation indices. $\kappa_{i}$ are couplings of mass dimension 1, and $\lambda$, $\lambda^{\prime}$, and $\lambda^{\prime\prime}$ are dimensionless couplings. Here, the factor $\frac{1}{2}$ is attached to the operators $LL\bar{E}$ and $\bar{U}\bar{D}\bar{D}$ because these operators possess anti-symmetry property such that $\lambda_{ijk}=-\lambda_{jik}$ and $\lambda^{\prime\prime}_{ijk}=-\lambda^{\prime\prime}_{ikj}$. In this work, we assume that only $\lambda$ and $\lambda^{\prime}$ couplings with certain flavor combinations are non-zero; this can be theoretically realized, as the bilinear term can be rotated away at a given energy scale Hall:1983id ; Dreiner:2003hw and the BNV term can be vanishing if, for example, the baryon-triality symmetry is imposed Dreiner:2005rd ; Ibanez:1991hv ; Martin:1997ns ; Grossman:1998py ; Dreiner:2006xw ; Dreiner:2007uj ; Bernhardt:2008jz .
@@
As discussed in Sec. 1, a light MeV-scale bino is allowed and can be long-lived for tiny RPV couplings. For the SHiNESS experiment, we focus on LLPs produced in $\pi^{+}$ and $\mu^{+}$ decays at rest and on the final-state signature of an electron-positron pair from the $\tilde{\chi}^{0}_{1}$ decays. The $\pi^{+}$ decays into a light bino can be induced by the $\lambda^{\prime}_{111}$ or $\lambda^{\prime}_{211}$ couplings, while the signal decays of the $\mu^{+}$ lepton and the bino can be mediated by various $\lambda$ couplings. For the light bino produced with the $\lambda^{\prime}_{211}$ coupling in $\pi^{+}\to\mu^{+}\tilde{\chi}^{0}_{1}$ decays, SHiNESS has no sensitivity, as we will explain later. In Table 1 we list four RPV benchmark scenarios we choose to study. In each of them, we assume only two RPV LNV couplings are non-vanishing, mediating production and decay of the lightest neutralino, respectively, except in one benchmark B2 where only a single coupling $\lambda_{121}$ is assumed to be non-zero and it leads to both the production and decay of the light bino. Specifically, in Table 1, we list the production and decay couplings for the light bino, as well as the corresponding signal processes. In addition, in some benchmark scenarios the production coupling can induce decays of the light bino in the interested mass range of the listed benchmark scenarios. For instance, in B1, the decay $\tilde{\chi}^{0}_{1}\to\pi^{0}\overset{(-)}{\nu}_{e}$ can, in principle, take place via $\lambda^{\prime}_{111}$, but the allowed bino mass window, $m_{\pi^{0}}<m_{\tilde{\chi}^{0}_{1}}<m_{\pi^{+}}$, is tiny, and therefore we ignore this case. Also, in B2, by definition the production and decay couplings are identical. Finally, in B3, the production coupling $\lambda_{121}$ can induce the $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ decays which are taken into account in the numerical computation.
@@
B1 $\lambda^{\prime}_{111}$ $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
B2 $\lambda_{121}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$ $\lambda_{121}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ Redundant Redundant
B3 $\lambda_{121}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$ $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$
B4 $\lambda_{312}$ $\mu^{+}\to\tilde{\chi}^{0}_{1}\nu_{\tau}e^{+}$ $\lambda_{131}$ $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\tau}e^{+}e^{-}$- --
@@
Table 1: Summary of the representative benchmark scenarios of the light binos in the RPV-SUSY that we study. For benchmarks B2-B4, the $\mu^{+}$ leptons are produced in $\pi^{+}\to\mu^{+}\nu_{\mu}$ decays at rest at the ESS and they also decay immediately at rest. The lightest neutralino is of Majorana nature and can hence decay to both charge-conjugated channels. “--” means either theoretically irrelevant, or kinematically forbidden.
@@
To compute the production rates of the light bino from pion and muon decays via $\lambda^{\prime}$ and $\lambda$ couplings, respectively, we resort to the analytic expressions provided in Refs. deVries:2015mfw ; Dreiner:2023gir , while the decay widths’ computation of the light bino via $\lambda$ couplings follows the procedure given in Ref. Dreiner:2023gir .
@@
We now discuss the present bounds on the relevant RPV couplings. These bounds stem from either measurements on low-energy observables Allanach:1999ic ; Dreiner:2023gir or recast of past searches for long-lived HNLs Dreiner:2023gir .
@@
For the coupling $\lambda^{\prime}_{111}$, both measurements of neutrinoless double-beta ($0\nu\beta\beta$) decays and existing bounds on long-lived HNLs provide the leading limits on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the lightest neutralino’s mass $m_{\tilde{\chi}^{0}_{1}}$. The bound from $0\nu\beta\beta$ decays on the coupling was obtained in Refs. Mohapatra:1986su ; Hirsch:1995zi ; Hirsch:1995ek ; Babu:1995vh . The reinterpretation of the HNL searches’ results in terms of the light binos in the RPV-SUSY was performed in Ref. Dreiner:2023gir , and its Fig. 2 shows that these bounds are stronger than those from the $0\nu\beta\beta$ decays for $m_{\tilde{\chi}^{0}_{1}}$ roughly between 30 MeV and 130 MeV.
The involved $\lambda$ couplings all receive bounds from low-energy processes, as listed in Ref. Allanach:1999ic . The bound on $\lambda_{121}$ originates from charged-current universality (C. C. universality) Barger:1989rk . The upper limit on $\lambda_{131}$ was derived from measurements of $R_{\tau}=\Gamma(\tau\to e\nu\bar{\nu})/\Gamma(\tau\to\mu\nu\bar{\nu})$ Barger:1989rk . Moreover, the current bound on $\lambda_{312}$ stems from measurements on $R_{\tau\mu}=\Gamma(\tau\to\mu\nu\bar{\nu})/\Gamma(\mu\to e\nu\bar{\nu})$ Ledroit:1998 .
@@
In addition, the $\lambda_{121}$ coupling is bounded from uncertainty on the muon decay width of $\mu^{\pm}\to e^{\pm}+$invisible ParticleDataGroup:2022pth ($\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$) as functions of $m_{\tilde{\chi}^{0}_{1}}$, as given in Fig. 6 of Ref. Dreiner:2023gir . The results are comparable with the bounds obtained from low-energy observables.
Coupling Relevant process Resulting upper bound References
@@
$\lambda^{\prime}_{111}$ $0\nu\beta\beta$ $5.2\cdot 10^{-4}\cdot\Big{(}\frac{m_{\tilde{e}}}{100\text{ GeV}}\Big{)}^{2}\cdot\sqrt{\frac{m_{\tilde{\chi}^{0}_{1}}}{100\text{ GeV}}}$ Mohapatra:1986su ; Hirsch:1995zi ; Hirsch:1995ek ; Babu:1995vh
HNL search recast a function of $m_{\tilde{\chi}^{0}_{1}}$ and $m_{\tilde{f}}$ Dreiner:2023gir
$\lambda_{121}$ C. C. universality $0.049\cdot\frac{m_{\tilde{e}_{R}}}{100\text{ GeV}}$ Barger:1989rk
$\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$ a function of $m_{\tilde{\chi}^{0}_{1}}$ and $m_{\tilde{f}}$ Dreiner:2023gir
$\lambda_{131}$ $R_{\tau}$ $0.062\cdot\frac{m_{\tilde{e}_{R}}}{100\text{ GeV}}$ Barger:1989rk
$\lambda_{312}$ $R_{\tau\mu}$ $0.062\cdot\frac{m_{\tilde{\mu}_{R}}}{100\text{ GeV}}$ Ledroit:1998
Table 2: Summary of the current constraints on the RPV couplings involved in the considered benchmark scenarios listed in Table 1. The relevant processes, resulting bounds, as well as the references where these bounds were obtained, are tabulated. For the $0\nu\beta\beta$ bound on $\lambda^{\prime}_{111},m_{\tilde{e}}$ labels the selectron’s mass. For the bounds on the $\lambda$ couplings, $\tilde{e}_{R}$ and $\tilde{\mu}_{R}$ denote right-handed selectron and smuon, respectively. The bound on $\lambda^{\prime}_{111}$ from HNL search recast and that on $\lambda_{121}$ from $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$ were derived in Ref. Dreiner:2023gir and given in the $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ and $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ planes, respectively, with no analytic expression.
In Table 2, we summarize the present bounds on these RPV couplings, including the relevant processes, resulting constraints, as well as the corresponding references.
@@
In this work, we assume degenerate sfermion masses $m_{\tilde{f}}$ for simplicity.
Finally, it is worth noting that most of these existing bounds on the RPV couplings scale with $m_{\tilde{f}}$. On the other hand, for the signal decay processes of the pion, muon, and light bino considered in this work, the decay widths are proportional to $\lambda^{(^{\prime})2}/m^{4}_{\tilde{f}}$ deVries:2015mfw ; Dreiner:2023gir . Therefore, these present bounds, when compared with the sensitivity reach of the SHiNESS experiment to this theoretical scenario as we will show later in Sec. 4, will be displayed for different sfermion masses.